OpenStudy (anonymous):

Can someone explain this question to me, please? Caroline runs 3 1/2 miles in 22 1/4 minutes. How many miles does she run per minute?

OpenStudy (math&ing001):

\[\frac{ 3+\frac{ 1 }{ 2 } }{ 22+\frac{ 1 }{ 4 } }=\frac{ x }{ 1 }\] Solve for x.

OpenStudy (math&ing001):

you're welcome :)

OpenStudy (anonymous):

^^ tankiez